Calculate EPS reported before stock split and stock dividend - During the fiscal year ended September 30, 2017, Worrell, Inc., had a 2-for-1 stock split and a 5% stock dividend. In its annual report for 2017, the company reported earnings per share for the year ended September 30, 2016, on a restated basis, of $1.20

Required: Calculate the originally reported earnings per share for the year ended September 30, 2016.
Answer #1

Solution:

Let share outstanding for the year ended Sep 30, 2016 = 100

Therefore computation of restated EPS, shares considered = 100 + stock split + Stock dividend

= 100 + 100 + (200*5%) = 210 share

It means nos of shares increased by 110%

Therefore Original reported earning per share for 2016 = $1.20 + $1.20 * 110% = $2.52